# High-low analysis and small cap decoupling over non-Archimedean local fields

### v1

@@ 2.4 Reduction to local estimates
$\begin{split}\mathfrak{a}^{p}\mu\big{(}\{x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\}\big{)}&\leq C2^{\frac{p}{2}-3}e^{2p-4}N^{p+E_{2}-4}(\log R)\sum_{0\leq k\leq N}(\#\mathcal{T}_{k}(\mathfrak{s}^{k}))^{p-4}\\ &\times\sum_{\tau_{k}\in\mathcal{T}_{k}(\mathfrak{s}^{k})}\left(\mathfrak{s}^{k}\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\right)^{\frac{p}{2}-1}\sum_{\gamma\prec\tau_{k}}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{r})}\|_{p}^{p},\end{split}$
@@
$\begin{split}\mathfrak{a}^{p}&\mu\big{(}\{x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\}\big{)}\\ &\leq C2^{\frac{p}{2}-3}e^{2p-4}(N+1)N^{p+E_{2}-4}(\log R)(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-4}\sum_{\tau_{k_{*}}\in\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})}\left(\mathfrak{s}^{k_{*}}\max_{\gamma_{k_{*}}\prec\tau_{k_{*}}}\#(\gamma\prec\gamma_{k_{*}})\right)^{\frac{p}{2}-1}\sum_{\gamma\prec\tau_{k_{*}}}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{r})}\|_{p}^{p}\\ &\leq C2^{\frac{p}{2}-3}e^{2p-4}(N+1)N^{p+E_{2}-4}(\log R)(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-3}\left(\mathfrak{s}^{k_{*}}\max_{\gamma_{k_{*}}\prec\tau_{k_{*}}}\#(\gamma\prec\gamma_{k_{*}})\right)^{\frac{p}{2}-1}\mathfrak{s}^{k_{*}}\lambda^{p}\mathfrak{r}\cdot R^{1+\beta}.\end{split}$
We claim that this amounts to a decoupling inequality for the $\lambda g^{(\lambda,\mathfrak{r})}$. That is,
@@
For each $\lambda,\mathfrak{r}$, and $\mathfrak{a}$, and each $p\geq 4$, we have the estimate
$\mathfrak{a}^{p}\mu\big{(}\{x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\}\big{)}\leq C2^{\frac{p}{2}-3}e^{4p-3}(N+1)N^{p+E_{2}-4}(\log R)\max\big{(}R^{\beta(p-\frac{p}{q}-1)-1}R^{\beta(\frac{p}{2}-\frac{p}{q})}\big{)}\Big{(}\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{r})}\|_{p}^{q}\Big{)}^{\frac{p}{q}}.$ (3.6)
@@
$\begin{split}\alpha^{p}\mu\big{(}\{x:|f(x)|>\alpha\}\big{)}&\leq\alpha^{p}\sum_{N^{-2}R^{-1/2}<\lambda\leq 1}\mu\Big{(}\Big{\{}x:|\lambda g^{(\lambda)}(x)|\geq\frac{\alpha}{\mathcal{Z}}\Big{\}}\Big{)}\\ &\leq\mathcal{Z}^{p+1}\Big{(}\frac{\alpha}{\mathcal{Z}}\Big{)}^{p}\mu\Big{(}\Big{\{}x:|\lambda_{*}g^{(\lambda_{*})}(x)|\geq\frac{\alpha}{\mathcal{Z}}\Big{\}}\Big{)},\end{split}$
(where we have abbreviated $\mathcal{Z}=2\log N+\frac{1}{2}\log R$) and hence, for a suitable $\mathfrak{r}$,
$\alpha^{p}\mu\big{(}\{x:|f(x)|>\alpha\}\big{)}\leq\left[\mathcal{Z}\log R\right]^{p+1}\Big{(}\frac{\alpha}{\mathcal{Z}\log R}\Big{)}^{p}\mu\Big{(}\Big{\{}x:|\lambda_{*}g^{(\lambda_{*},\mathfrak{r})}(x)|\geq\frac{\alpha}{\mathcal{Z}\log R}\Big{\}}\Big{)},$
@@
$\begin{split}\alpha^{p}\mu\big{(}\{x:|f(x)|>\alpha\}\big{)}|&\leq C2^{\frac{p}{2}-3}e^{4p-3}\mathcal{Z}^{p+1}(N+1)N^{p+E_{2}-4}(\log R)^{p+2}\\ &\times\max\big{(}R^{\beta(p-\frac{p}{q}-1)-1}R^{\beta(\frac{p}{2}-\frac{p}{q})}\big{)}\Big{(}\sum_{\gamma}\|\lambda_{*}g_{\gamma}^{(\lambda_{*},\mathfrak{r})}\|_{p}^{q}\Big{)}^{\frac{p}{q}}.\end{split}$
Finally, we note that each $\lambda_{*}g_{\gamma}^{(\lambda_{*},\mathfrak{r})}$ is obtained by taking a subsum of a partition of unity applied to $f_{\gamma}$, so we conclude that
$\begin{split}\alpha^{p}\mu\big{(}\{x:|f(x)|>\alpha\}\big{)}|&\leq C2^{\frac{p}{2}-3}e^{4p-3}\mathcal{Z}^{p+1}(N+1)N^{p+E_{2}-4}(\log R)^{p+2}\\ &\times\max\big{(}R^{\beta(p-\frac{p}{q}-1)-1}R^{\beta(\frac{p}{2}-\frac{p}{q})}\big{)}\Big{(}\sum_{\gamma}\|f_{\gamma}\|_{p}^{q}\Big{)}^{\frac{p}{q}},\end{split}$ (3.7)
@@
$R^{-\frac{p}{2}+2}\leq R^{\frac{p}{2}+1-\beta(p-1)}\leq\max_{\gamma}\|f_{\gamma}\|_{p}^{p},$ (3.8)
[…]
@@
$D_{p_{\beta},q_{\beta}}^{\mathbb{K}}(R;\beta)\leq 10^{6}(\log\mathbf{p})^{-16}\mathbf{p}^{12}(\log R)^{17+6\beta^{-1}}\max\Big{(}R^{\beta(\frac{p_{\beta}}{2}-\frac{p_{\beta}}{q_{\beta}})},R^{\beta(p_{\beta}-\frac{p_{\beta}}{q_{\beta}}-1)-1}\Big{)}\Big{(}\sum_{\gamma}\|f_{\gamma}\|_{p_{\beta}}^{q_{\beta}}\Big{)}^{\frac{p_{\beta}}{q_{\beta}}}.$
Remark 3.4.
The condition $R\geq\mathbf{p}^{32}$, equivalent to $N\geq 16$, is arranged to make the universal constant at most $10^{6}$. Asymptotically, one can arrange for the constant to be $\leq 4885$.
@@
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 2^{-5}e^{4}(\log\mathbf{p})^{-8}\mathbf{p}^{5}(\log R)^{8}R^{\beta}.$
The latter estimate is established in Prop. (3.5) below; there a different small cap decoupling argument is used than the one expounded so far, which is more efficient in the subcritical regime $2\leq p\leq 4$.
For each $3\leq p\leq\infty$, write $\frac{1}{q_{p}}=1-\frac{3}{p}$. Write $p_{\beta}=2+\frac{2}{\beta}$ and $q_{\beta}=q_{p_{\beta}}$. Let $\frac{3}{p}+\frac{1}{q}\leq 1$. Suppose that $2+\frac{2}{\beta}\leq p$ and $\beta<1$. Then, from the two inequalities
$D_{p_{\beta},q_{\beta}}^{\mathbb{K}}(R;\beta)\leq 10^{6}(\log\mathbf{p})^{-16}\mathbf{p}^{12}(\log R)^{17+6\beta^{-1}}R^{\beta(\frac{p_{\beta}}{2}-\frac{p_{\beta}}{q_{\beta}})},$
@@
$D_{p,q_{p}}^{\mathbb{K}}(R;\beta)\leq 10^{6}(\log\mathbf{p})^{-16}\mathbf{p}^{12}(\log R)^{17+6\beta^{-1}}R^{\beta(p-\frac{p}{q_{p}}-1)-1}.$
@@
$D_{p,q}^{\mathbb{K}}(R;\beta)\leq 10^{6}(\log\mathbf{p})^{-16}\mathbf{p}^{12}(\log R)^{17+6\beta^{-1}}R^{\beta(p-\frac{p}{q}-1)-1}$
Next, suppose that $4\leq p\leq 2+\frac{2}{\beta}$. By Lemma (3.5), we may assume $\beta<1$. Then, from the inequalities
$D_{p_{\beta},q_{\beta}}^{\mathbb{K}}(R;\beta)\leq 10^{6}(\log\mathbf{p})^{-16}\mathbf{p}^{12}(\log R)^{17+6\beta^{-1}},$
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 2^{-5}e^{4}(\log\mathbf{p})^{-8}\mathbf{p}^{5}(\log R)^{8}R^{\beta}$
@@
$D_{p,q_{p}}^{\mathbb{K}}(R;\beta)\leq 10^{6}(\log\mathbf{p})^{-8}\mathbf{p}^{5+\frac{7\beta}{2-2\beta}(p-4)}(\log R)^{8+\frac{9\beta+6}{2-2\beta}(p-4)}R^{\beta(\frac{p}{2}-\frac{p}{q_{p}})}.$
@@
$D_{p,q}^{\mathbb{K}}(R;\beta)\leq 10^{6}(\log\mathbf{p})^{-8}\mathbf{p}^{5+\frac{7\beta}{2-2\beta}(p-4)}(\log R)^{8+\frac{3\beta+2}{1-\beta}(p-4)}R^{\beta(\frac{p}{2}-\frac{p}{q})}.$
It remains to consider the regime $3\leq p\leq 4$. Then, from the estimates
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 2(\log\mathbf{p})^{-8}\mathbf{p}^{5}(\log R)^{8}R^{\beta},$
@@
$D_{3,3}^{\mathbb{K}}(R;\beta)\leq 2(\log\mathbf{p})^{-4}\mathbf{p}^{5/2}(\log R)^{4}R^{\frac{\beta}{2}},$
@@
$D_{3,\infty}^{\mathbb{K}}(R;\beta)\leq 2(\log\mathbf{p})^{-4}\mathbf{p}^{5/2}(\log R)^{4}R^{\frac{3\beta}{2}}.$
@@
$D_{p,q}^{\mathbb{K}}(R;\beta)\leq 2(\log\mathbf{p})^{-4(p-2)}\mathbf{p}^{5/2(p-2)}(\log R)^{4(p-2)}R^{\beta(\frac{p}{2}-\frac{p}{q})}.$
To close the argument, it remains only to prove Lemmas (3.3) and (3.5).
@@
The proof is by casework on $p$ and $k$. Suppose $R_{k}\geq R^{1-\beta}$. By a straightforward calculation (see Case 2 of the proof of Theorem 5 in [16]),
(#$(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-4}\left(\mathfrak{s}^{k_{*}}\max_{\begin{subarray}{c}\gamma_{k_{*}}\prec\tau_{k_{*}}\\ \tau_{k_{*}}\in\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\end{subarray}}\#(\gamma\prec\gamma_{k_{*}})\right)^{\frac{p}{2}-1}\leq R^{\beta(p-\frac{p}{q}-1)-1}\big{(}\mathfrak{s}^{k_{*}}\times\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\big{)}^{\frac{p}{q}-1}.$
@@
$\begin{split}\mathfrak{a}^{p}&\mu\big{(}\big{\{}x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\big{\}}\big{)}\\ &\leq C2^{\frac{p}{2}-3}e^{2p-4}(N+1)N^{p+E_{2}-4}(\log R)R^{\beta(p-\frac{p}{q}-1)-1}\big{(}\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\big{)}^{\frac{p}{q}}(\mathfrak{s}^{k_{*}})^{\frac{p}{q}}\lambda^{p}\mathfrak{r}R^{1+\beta}\\ &\leq C2^{\frac{p}{2}-3}e^{4p-3}(N+1)N^{p+E_{2}-4}(\log R)R^{\beta(p-\frac{p}{q}-1)-1}\Big{(}\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{r})}\|_{p}^{q}\Big{)}^{\frac{p}{q}}.\end{split}$ (3.10)
Suppose instead that $R_{k}\leq R^{1-\beta}$, and $2+\frac{2}{\beta}\leq p\leq 6$. Then, from the inequality
$1\leq R^{\beta(\frac{p}{2}-1)}R^{-1},$
@@
[#$[\#(\gamma\prec\gamma_{k})]^{\frac{p}{2}-1}\leq R^{-1}R^{\beta(\frac{p}{2}-1)}R_{k}^{3-\frac{p}{2}}.$
@@
$R^{\beta(p-\frac{p}{q}-3)}[\#(\gamma\prec\gamma_{k})]^{\frac{p}{2}-1}\leq R^{\beta(p-\frac{p}{q}-1)-1}\Big{(}R^{-\beta}R_{k}\Big{)}^{3-\frac{p}{2}},$
which implies (using $\frac{1}{q}+\frac{3}{p}\leq 1$)
$(\mathfrak{s}^{k_{*}}\times\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-3}[\#(\gamma\prec\gamma_{k})]^{\frac{p}{2}-1}\leq R^{\beta(p-\frac{p}{q}-1)-1}\Big{(}R^{-\beta}R_{k}\Big{)}^{3-\frac{p}{2}}(\#\gamma)^{\frac{p}{q}}.$
@@
(#$(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-4}(\#\gamma\prec\tau_{k})^{\frac{p}{2}-1}\leq(\mathfrak{s}^{k_{*}})^{3-\frac{p}{2}}(\mathfrak{s}^{k_{*}}\times\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-4},$
@@
$\mathfrak{s}^{k_{*}}(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{p-3}(\#\gamma\prec\tau_{k})^{\frac{p}{2}-1}[\#(\gamma\prec\gamma_{k})]^{\frac{p}{2}-1}\leq R^{\beta(p-\frac{p}{q}-1)-1}(\#\gamma)^{\frac{p}{q}}.$
@@
$\begin{split}\mathfrak{a}^{p}\mu\big{(}\{x:|\lambda g^{(\lambda,\mathfrak{r})}(x)|>\mathfrak{a}\}\big{)}&\leq C2^{\frac{p}{2}-3}e^{2p-4}(N+1)N^{p+E_{2}-4}(\log R)R^{\beta(p-\frac{p}{q}-1)-1}(\#\gamma)^{\frac{p}{q}}\lambda^{p}\mathfrak{r}\cdot R^{1+\beta}\\ &\leq C2^{\frac{p}{2}-3}e^{4p-3}(N+1)N^{p+E_{2}-4}(\log R)R^{\beta(p-\frac{p}{q}-1)-1}\Big{(}\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{r})}\|_{p}^{q}\Big{)}^{\frac{p}{q}}.\end{split}$
It remains to consider the case $R_{k}\leq R^{1-\beta}$, $4\leq p\leq 2+\frac{2}{\beta}$. In this case, from the inequalities
$R_{k}^{\frac{p}{2}-3}\leq 1,\quad\mathfrak{s}^{k_{*}}\leq RR_{k}^{-1},\quad\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\leq R_{k},$

### v2

@@ 2.4 Reduction to local estimates
$\begin{split}\mathfrak{a}^{r}\mu\big(\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\}\big)&\leq C2^{\frac{r}{2}-3}e^{2r-4}N^{r+E_{2}-4}(\log R)\sum_{0\leq k\leq N}(\#\mathcal{T}_{k}(\mathfrak{s}^{k}))^{r-4}\\ &\times\sum_{\tau_{k}\in\mathcal{T}_{k}(\mathfrak{s}^{k})}\left(\mathfrak{s}^{k}\max_{\gamma_{k}\prec\tau_{k}}\#(\gamma\prec\gamma_{k})\right)^{\frac{r}{2}-1}\sum_{\gamma\prec\tau_{k}}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{t})}\|_{r}^{r},\end{split}$
@@
$\begin{split}\mathfrak{a}^{p}&\mu\big(\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\}\big)\\ &\leq C2^{\frac{r}{2}-3}e^{2r-4}(N+1)N^{r+E_{2}-4}(\log R)(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-4}\sum_{\tau_{k_{*}}\in\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})}\left(\mathfrak{s}^{k_{*}}\max_{\gamma_{k_{*}}\prec\tau_{k_{*}}}\#(\gamma\prec\gamma_{k_{*}})\right)^{\frac{r}{2}-1}\sum_{\gamma\prec\tau_{k_{*}}}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{t})}\|_{r}^{r}\\ &\leq C2^{\frac{r}{2}-3}e^{2r-4}(N+1)N^{r+E_{2}-4}(\log R)(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-3}\left(\mathfrak{s}^{k_{*}}\max_{\gamma_{k_{*}}\prec\tau_{k_{*}}}\#(\gamma\prec\gamma_{k_{*}})\right)^{\frac{r}{2}-1}\mathfrak{s}^{k_{*}}\lambda^{r}\mathfrak{t}\cdot R^{1+\beta}.\end{split}$
We claim that this amounts to a decoupling inequality for the $\lambda g^{(\lambda,\mathfrak{t})}$. That is,
@@
For each $\lambda,\mathfrak{t}$, and $\mathfrak{a}$, and each $r\geq 4$, we have the estimate
$\mathfrak{a}^{r}\mu\big(\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\}\big)\leq C2^{\frac{r}{2}-3}e^{4r-3}(N+1)N^{r+E_{2}-4}(\log R)\max\big(R^{\beta(r-\frac{r}{q}-1)-1},R^{\beta(\frac{r}{2}-\frac{r}{q})}\big)\Big(\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{t})}\|_{r}^{q}\Big)^{\frac{r}{q}}.$ (3.6)
@@
$\begin{split}\alpha^{r}\mu\big(\{x:|f(x)|>\alpha\}\big)&\leq\alpha^{r}\sum_{N^{-2}R^{-1/2}<\lambda\leq 1}\mu\Big(\Big\{x:|\lambda g^{(\lambda)}(x)|\geq\frac{\alpha}{\mathcal{Z}}\Big\}\Big)\\ &\leq\mathcal{Z}^{r+1}\Big(\frac{\alpha}{\mathcal{Z}}\Big)^{r}\mu\Big(\Big\{x:|\lambda_{*}g^{(\lambda_{*})}(x)|\geq\frac{\alpha}{\mathcal{Z}}\Big\}\Big),\end{split}$
(where we have abbreviated $\mathcal{Z}=2\log N+\frac{1}{2}\log R+1$) and hence, for a suitable $\mathfrak{t}$,
$\alpha^{r}\mu\big(\{x:|f(x)|>\alpha\}\big)\leq\left[\mathcal{Z}\log R\right]^{r+1}\Big(\frac{\alpha}{\mathcal{Z}\log R}\Big)^{r}\mu\Big(\Big\{x:|\lambda_{*}g^{(\lambda_{*},\mathfrak{t})}(x)|\geq\frac{\alpha}{\mathcal{Z}\log R}\Big\}\Big),$
@@
$\begin{split}\alpha^{r}\mu\big(\{x:|f(x)|>\alpha\}\big)|&\leq C2^{\frac{r}{2}-3}e^{4r-3}\mathcal{Z}^{r+1}(N+1)N^{r+E_{2}-4}(\log R)^{r+2}\\ &\times\max\big(R^{\beta(r-\frac{r}{q}-1)-1},R^{\beta(\frac{r}{2}-\frac{r}{q})}\big)\Big(\sum_{\gamma}\|\lambda_{*}g_{\gamma}^{(\lambda_{*},\mathfrak{t})}\|_{r}^{q}\Big)^{\frac{r}{q}}.\end{split}$
Finally, we note that each $\lambda_{*}g_{\gamma}^{(\lambda_{*},\mathfrak{t})}$ is obtained by taking a subsum of a partition of unity applied to $f_{\gamma}$, so we conclude that
$\begin{split}\alpha^{r}\mu\big(\{x:|f(x)|>\alpha\}\big)|&\leq C2^{\frac{r}{2}-3}e^{4r-3}\mathcal{Z}^{r+1}(N+1)N^{r+E_{2}-4}(\log R)^{r+2}\\ &\times\max\big(R^{\beta(r-\frac{r}{q}-1)-1},R^{\beta(\frac{r}{2}-\frac{r}{q})}\big)\Big(\sum_{\gamma}\|f_{\gamma}\|_{r}^{q}\Big)^{\frac{r}{q}},\end{split}$ (3.7)
@@
$R^{-\frac{r}{2}+2}\leq R^{\frac{r}{2}+1-\beta(r-1)}\leq\max_{\gamma}\|f_{\gamma}\|_{r}^{r},$ (3.8)
$\int_{R^{-1/2}}^{R^{1/2}}r\alpha^{r-1}\mu(U_{\alpha})d\alpha\leq r(\log R)\sup_{\alpha\in[R^{-1/2},R^{1/2}]}\alpha^{r}\mu(U_{\alpha}),$ (3.9)
@@
$D_{r_{\beta},q_{\beta}}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}R^{\beta(\frac{r_{\beta}}{2}-\frac{r_{\beta}}{q_{\beta}})}.$
@@
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 2R^{\beta}.$
The latter estimate follows trivially from Córdoba–Fefferman and flat decoupling.
For each $3\leq r\leq\infty$, write $\frac{1}{q_{r}}=1-\frac{3}{r}$. Write $r_{\beta}=2+\frac{2}{\beta}$ and $q_{\beta}=q_{r_{\beta}}$. Let $\frac{3}{r}+\frac{1}{q}\leq 1$. Suppose that $2+\frac{2}{\beta}\leq r$ and $\beta<1$. Then, from the two inequalities
$D_{p_{\beta},q_{\beta}}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}R^{\beta(\frac{r_{\beta}}{2}-\frac{r_{\beta}}{q_{\beta}})},$
@@
$D_{r,q_{r}}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}R^{\beta(r-\frac{r}{q_{r}}-1)-1}.$
@@
$D_{r,q}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}R^{\beta(r-\frac{r}{q}-1)-1}.$
Next, suppose that $4<r\leq 2+\frac{2}{\beta}$. We may assume $\beta<1$. Then, from the inequalities
$D_{r_{\beta},q_{\beta}}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}R^{\beta(\frac{r_{\beta}}{2}-\frac{r_{\beta}}{q_{\beta}})},$
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 2R^{\beta}$
@@
$D_{r,q_{r}}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}R^{\beta(\frac{r}{2}-\frac{r}{q_{r}})}.$
@@
$D_{r,q}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}R^{\beta(\frac{r}{2}-\frac{r}{q})}.$
It remains to consider the regime $3\leq r\leq 4$. Then, from the estimates
$D_{4,4}^{\mathbb{K}}(R;\beta)\leq 2R^{\beta},$
@@
$D_{3,3}^{\mathbb{K}}(R;\beta)\leq\sqrt{2}R^{\frac{\beta}{2}},$
@@
$D_{3,\infty}^{\mathbb{K}}(R;\beta)\leq\sqrt{2}R^{\frac{3\beta}{2}}.$
@@
$D_{r,q}^{\mathbb{K}}(R;\beta)\leq 2R^{\beta(\frac{r}{2}-\frac{r}{q})}.$
In all, we conclude the estimate
$D_{r,q}^{\mathbb{K}}(R;\beta)\leq\frac{e^{5+2\beta^{-1}}}{90\cdot 2^{3\beta^{-1}}}\frac{\mathbf{q}^{-3}\mathbf{p}^{10}}{(\log\mathbf{p})^{9+2\beta^{-1}}}(\log R)^{17+6\beta^{-1}}\max\Big(R^{\beta(r-\frac{r}{q}-1)-1},R^{\beta(\frac{r}{2}-\frac{r}{q})}\Big).$
The result follows.
To close the argument, it remains only to prove Lemma (3.3).
@@
The proof is by casework on $r$ and $k$. Suppose $R_{k}\geq R^{1-\beta}$. By a straightforward calculation (see Case 2 of the proof of Theorem 5 in [17]),
$(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-4}\left(\mathfrak{s}^{k_{*}}\max_{\begin{subarray}{c}\gamma_{k_{*}}\prec\tau_{k_{*}}\\ \tau_{k_{*}}\in\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\end{subarray}}\#(\gamma\prec\gamma_{k_{*}})\right)^{\frac{r}{2}-1}\leq R^{\beta(r-\frac{r}{q}-1)-1}\big(\mathfrak{s}^{k_{*}}\times\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\big)^{\frac{r}{q}-1}.$
@@
$\begin{split}\mathfrak{a}^{r}&\mu\big(\big\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\big\}\big)\\ &\leq C2^{\frac{r}{2}-3}e^{2r-4}(N+1)N^{r+E_{2}-4}(\log R)R^{\beta(r-\frac{r}{q}-1)-1}\big(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\big)^{\frac{r}{q}}(\mathfrak{s}^{k_{*}})^{\frac{r}{q}}\lambda^{r}\mathfrak{t}R^{1+\beta}\\ &\leq C2^{\frac{r}{2}-3}e^{4r-3}(N+1)N^{r+E_{2}-4}(\log R)R^{\beta(r-\frac{r}{q}-1)-1}\Big(\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{t})}\|_{r}^{q}\Big)^{\frac{r}{q}}.\end{split}$ (3.10)
Suppose instead that $R_{k}\leq R^{1-\beta}$, and $2+\frac{2}{\beta}\leq r\leq 6$. Then, from the inequality
$1\leq R^{\beta(\frac{r}{2}-1)}R^{-1},$
@@
$(\gamma\prec\gamma_{k})]^{\frac{r}{2}-1}\leq R^{-1}R^{\beta(\frac{r}{2}-1)}R_{k}^{3-\frac{r}{2}}.$
@@
$R^{\beta(r-\frac{r}{q}-3)}[\#(\gamma\prec\gamma_{k})]^{\frac{r}{2}-1}\leq R^{\beta(r-\frac{r}{q}-1)-1}\Big(R^{-\beta}R_{k}\Big)^{3-\frac{r}{2}},$
which implies (using $\frac{1}{q}+\frac{3}{r}\leq 1$)
$(\mathfrak{s}^{k_{*}}\times\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-3}[\#(\gamma\prec\gamma_{k})]^{\frac{r}{2}-1}\leq R^{\beta(r-\frac{r}{q}-1)-1}\Big(R^{-\beta}R_{k}\Big)^{3-\frac{r}{2}}(\#\gamma)^{\frac{r}{q}}.$
@@
$(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-4}(\#\gamma\prec\tau_{k})^{\frac{r}{2}-1}\leq(\mathfrak{s}^{k_{*}})^{3-\frac{r}{2}}(\mathfrak{s}^{k_{*}}\times\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-4},$
@@
$\mathfrak{s}^{k_{*}}(\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}}))^{r-3}(\#\gamma\prec\tau_{k})^{\frac{r}{2}-1}[\#(\gamma\prec\gamma_{k})]^{\frac{r}{2}-1}\leq R^{\beta(r-\frac{r}{q}-1)-1}(\#\gamma)^{\frac{r}{q}}.$
@@
$\begin{split}\mathfrak{a}^{r}\mu\big(\{x:|\lambda g^{(\lambda,\mathfrak{t})}(x)|>\mathfrak{a}\}\big)&\leq C2^{\frac{r}{2}-3}e^{2r-4}(N+1)N^{r+E_{2}-4}(\log R)R^{\beta(r-\frac{r}{q}-1)-1}(\#\gamma)^{\frac{r}{q}}\lambda^{r}\mathfrak{t}\cdot R^{1+\beta}\\ &\leq C2^{\frac{r}{2}-3}e^{4r-3}(N+1)N^{r+E_{2}-4}(\log R)R^{\beta(r-\frac{r}{q}-1)-1}\Big(\sum_{\gamma}\|\lambda g_{\gamma}^{(\lambda,\mathfrak{t})}\|_{r}^{q}\Big)^{\frac{r}{q}}.\end{split}$
It remains to consider the case $R_{k}\leq R^{1-\beta}$, $4\leq r\leq 2+\frac{2}{\beta}$. In this case, from the inequalities
$R_{k}^{\frac{r}{2}-3}\leq 1,\quad\mathfrak{s}^{k_{*}}\leq RR_{k}^{-1},\quad\#\mathcal{T}_{k_{*}}(\mathfrak{s}^{k_{*}})\leq R_{k},$